OpenStudy (anonymous):

A certain rectangles length is 1 feet longer than its width. if the area of the rectangle is 110 square ft, find its dimensions

OpenStudy (anonymous):

let width= x so, length=x+1 area= length x breadth = x(x+1)=110 or, x^2+x-110=0 x=10 or -11 since x represents dimension it cannot be negative.So, x=10 so width=10 and length=11. So it was 10 ft by 11 ft Hope this helps! :D